the third angle in a triangle whose other angles measure 35 and 66 degrees
Answer by richard1234(7193)

The angles in a triangle add up to 180. If we let the third angle be x, then

35 + 66 + x = 180

x = 180 - 35 - 66 = 79 (degrees)